Two dices are tossed together, what is the probability of having a multiple of 12.

1
Expert's answer
2022-04-11T16:01:36-0400

The only multiple that one can get is "12". The reason is that one can get "6" or less from one dice. It is possible to get "12" in case we get "6" from both dices. The probability is "\frac{1}{6}\cdot\frac{1}{6}=\frac{1}{36}" due to the multiplication rule of probability. Thus, the answer is "\frac{1}{36}".